A total of 56 different two item permutations can be created from a pile of clothes containing?
A. 9 garments
B. 8 garments
C. 10 garments
D. 20 garments

Answer by Edwin McCravy(20054)

A. 9P2 = (9)(8) = 72
B. 8P2 = (8)(7) = 56   <--the correct one
C. 10P2 = (10)(9) = 90
D. 20P2 = (20)(19) = 380 

Edwin